OpenStudy (anonymous):

5h + 2(11 - h) = -5 What is h?

OpenStudy (anonymous):

5h + 22 - 2h = -5 3h = -27 h = -9

OpenStudy (anonymous):

ohhh thankk youuu

OpenStudy (anonymous):

\[5h+2(11-h)=-5\\5h+22-2h=-5\\3h=-27\\h=-9\]

OpenStudy (anonymous):

-9